# Chiral Domain Walls Induced by Radially Magnetized Nanotube Geometry

### v1

@@ Abstract
We theoretically investigate chiral domain walls (DWs) formed in radially magnetized nanotubes with perpendicular magnetic anisotropy (PMA). Unlike tubes with an easy axis of magnetization oriented in other directions, radially magnetized tubes exhibit geometrical effects arising not only from exchange interactions but also from magnetostatic interactions, which lead to Dzyaloshinskii-Moriya interaction (DMI)-like effects. We derive expressions for the effective magnetic fields acting on DWs within PMA nanotubes and quantify SOT-driven DW motion using an analytical one-dimensional model, which is validated by micromagnetic simulations. Our results show that the DMI-like field due to magnetostatic interactions can be as significantly as the contribution of the material-induced DMI in nanotubes with diameters below 100 nm. This implies that the direction and speed of DW motion in PMA nanotubes could differ from those observed in nanoribbons composed of the same material. Furthermore, we demonstrate that DW velocity can be effectively controlled by adjusting the tube diameter and exchange stiffness constant of the magnetic layer, rather than relying solely on the material-induced DMI. These insights are expected to greatly expand the potential applications of PMA nanotube-based DW devices.
@@ I Introduction
In this study, we focus on nanotubes as a promising form for future 3D magnetic devices, alongside nanoribbons and nanocylinders. While fabricating 3D nanoribbon structures requires advanced techniques to deposit magnetic layers onto sidewalls [16], nanotubes and nanocylinders could be more readily fabricated using existing methods such as electroplating [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30] and atomic layer deposition [31, 32]. Moreover, nanotube structures can achieve a smaller cross-sectional area compared to non-hollow nanocylinders, thereby reducing the shift current for DWs.
@@
Previous theoretical research on the properties of DWs in nanotubes has primarily focused on systems with in-plane magnetized anisotropy (IMA) [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44]. These studies have predominantly considered DW motion driven by conventional spin-transfer torque (STT) and Ørsted field [37, 44]. In contrast, this study aims to establish a theoretical framework for DW dynamics in radially magnetized nanotubes induced by perpendicular magnetized anisotropy (PMA) materials [45]. Such systems are expected to utilize spin-orbit torque (SOT) [46, 47, 48], potentially enabling more efficient DW motion compared to traditional driving forces. Therefore, PMA nanotube-based DW shift registers are considered highly promising candidates for low power consumption data storage applications.
@@
Generally, geometrical effects are known to arise from exchange interactions [49, 50]; however, it has been predicted that geometrical effects due to magnetostatic interactions can also occur, depending on the curvature characteristics and the easy axis of magnetization [51]. Previous research indicates that such magnetostatic interactions induce geometrical effects only in systems with a radial easy axis of magnetization in tube structures, leading to energy contributions similar to Dzyaloshinskii-Moriya interaction (DMI) [52, 53]. This suggests the potential induction of chirality in DWs within tube structures. The magnitude of DMI is crucial for the efficiency of SOT-driven DW motion [46], making it essential to theoretically predict the impact of these geometrical effects, to evaluate the potential applications of PMA nanotubes.
@@
The purpose of this study is to elucidate the characteristics of chiral DWs in PMA nanotubes and to quantitatively predict the influence of geometrical effects on SOT-driven DW motion. Specifically, we derive expressions for the effective magnetic fields acting on DWs due to magnetostatic interactions and formulate an analytical model for SOT-driven DW motion. The validity of this model is confirmed through micromagnetic (µM) simulations.
@@
The remainder of this paper is organized as follows. In Section II, we describe the model of PMA nanotubes. Section III presents the expressions for effective magnetic fields due to geometrical effects in PMA nanotubes and discusses the characteristics of these fields and the stability of DWs as a function of nanotube diameter. In Section IV, we present the results of SOT-driven DW motion calculations, comparing the outcomes from the analytical one-dimensional model (1DM) [54, 55, 46, 56, 57] and µM simulations. We also discuss the comparative shift characteristics of PMA nanotubes and nanoribbons, as well as the impact of variations in tube diameter and magnetic properties on these characteristics. Finally, Section V summarizes the key findings and implications of this study.
@@ II Model of radially magnetized nanotube
Figure 1: Schematic illustration of a radially magnetized nanotube containing a DW. The magnetic structure for a topological charge of $Q=+1$ is presented. The enlarged diagram shows the magnetization configuration of the DW as seen from the inside of the tube.
[…]
For PMA layers of nanotubes, the effective magnetic anisotropy constant is approximately given by
$K_{{\rm eff}}=K_{{\rm u}}-2\pi M_{{\rm s}}^{2}-A/[R_{1}(R_{1}-t/2)],$ (1)
where the second term represents the demagnetization energy density, and the third term accounts for the exchange energy density associated with the radial domain state [70], which is specific to tube structures. Figure 2 shows the phase diagram of magnetic domain structures in the PMA nanotube, validated through µM simulations (see Appendix A). When $K_{{\rm eff}}>0$, the easy axes of magnetization are oriented radially; otherwise, the easy axis aligns along the axial direction.
Figure 2: Phase diagram of magnetic domain structures in the PMA nanotube. The dashed line represents the boundary of stable states as determined by Eq. (1). The markers indicate conditions validated through µM simulations. The following parameters are used; $A=1\,{\rm uerg/cm}$ and $t=1\,{\rm nm}$.
In PMA nanotubes, the energy densities associated with the PMA, DMI and exchange interactions can be expressed respectively as $\varepsilon_{{\rm PMA}}=-K_{{\rm eff}}m_{r}^{2}$, $\varepsilon_{{\rm DMI}}=D\left(m_{r}\frac{\partial}{\partial z}m_{z}-m_{z}\frac{\partial}{\partial z}m_{r}\right)$, and $\varepsilon_{{\rm exc}}=A\sum_{i=x,y,z}\left[\left(\frac{\partial m_{i}}{\partial r}\right)^{2}+\frac{1}{r^{2}}\left(\frac{\partial m_{i}}{\partial\varphi}\right)^{2}+\left(\frac{\partial m_{i}}{\partial z}\right)^{2}\right]$, where $m_{x}=m_{r}\cos\varphi-m_{\varphi}\sin\varphi$ and $m_{y}=m_{r}\sin\varphi+m_{\varphi}\cos\varphi$. The energy of magnetistatic interaction is expressed as:
@@
$E_{{\rm dem,bb}}$ $=\frac{1}{2}\int{\rm d}^{3}\bm{x}\int{\rm d}^{3}\bm{x}^{\prime}\frac{\rho_{{\rm m}}\left(\bm{x}\right)\rho_{{\rm m}}\left(\bm{x}^{\prime}\right)}{\left|\bm{x}-\bm{x}^{\prime}\right|},$
$E_{{\rm dem,ss}}$ $=\frac{1}{2}\int{\rm d}^{2}\bm{x}\int{\rm d}^{2}\bm{x}^{\prime}\frac{\sigma_{{\rm m}}\left(\bm{x}\right)\sigma_{{\rm m}}\left(\bm{x}^{\prime}\right)}{\left|\bm{x}-\bm{x}^{\prime}\right|},$
$E_{{\rm dem,bs}}$ $=\int{\rm d}^{3}\bm{x}\int{\rm d}^{2}\bm{x}^{\prime}\frac{\rho_{{\rm m}}\left(\bm{x}\right)\sigma_{{\rm m}}\left(\bm{x}^{\prime}\right)}{\left|\bm{x}-\bm{x}^{\prime}\right|},$
where $\sigma_{{\rm m}}\left(\bm{x}\right)=M_{\rm s}\bm{m}\left(\bm{x}\right)\cdot\bm{n}_{\perp}\left(\bm{x}\right)$ is the surface magnetic charge density, with $\bm{n}_{\perp}\left(\bm{x}\right)$ being the normal vector of the tube surface, and $\rho_{{\rm m}}\left(\bm{x}\right)=-M_{\rm s}\nabla\cdot\bm{m}\left(\bm{x}\right)$ is the volume magnetic charge density [70].
@@ III Characteristics of domain walls
In this section, we derive the energy densities per nanotube cross section and present expressions for the effective fields acting on the DWs in PMA nanotubes.
[…]
The PMA, DMI, and exchange energy densities per nanotube cross section are expressed as $w_{{\rm PMA,DMI,exc}}=\int{\rm d}z\,\varepsilon_{{\rm PMA,DMI,exc}}$. The integral result of the PMA term is $w_{{\rm PMA}}=2\sqrt{AK_{{\rm eff}}}$, which does not depend on $\phi$ and therefore does not affect the DW dynamics (see Eq. (9)).
@@
$w_{{\rm DMI}}\left(\phi\right)=-\pi\lambda M_{{\rm s}}H_{{\rm DMI}}Q\cos\phi,\$ (3)
where $H_{{\rm DMI}}=\frac{D}{M_{{\rm s}}\lambda}$ is the material-induced DMI field [71, 56, 72].
The integral result of the $\phi$-dependent exchange interaction term is given by
$w_{{\rm exc}}\left(\phi\right)=\lambda M_{{\rm s}}H_{{\rm a,exc}}^{{\rm tube}}\cos^{2}\phi,\$ (4)
where $H_{{\rm a,exc}}^{{\rm tube}}=-\frac{4A}{M_{{\rm s}}}\frac{1}{R_{1}^{2}-R_{0}^{2}}\ln\frac{R_{1}}{R_{0}}$ is the curvilinear-geometry-induced anisotropy-like field [49, 50]. This result is obtained by integrating $\varepsilon_{{\rm exc}}=\frac{A}{r^{2}}\left(1+\frac{r^{2}}{\lambda^{2}}{\rm sech}^{2}\frac{z-X}{\lambda}-{\rm sech}^{2}\frac{z-X}{\lambda}\cos^{2}\phi\right)$ with respect to $X$ and is consistent with the result for IMA tubes with an azimuthal easy axis of magnetization [44]. In the case where $H_{{\rm a,exc}}^{{\rm tube}}$ is positive (negative), the Bloch (Néel)-type DW is favored.
The energy density of the $\phi$-dependent magnetostatic interaction term is given by the following equation, as derived in Appendix B:
$w_{{\rm dem}}\left(\phi\right)=\lambda M_{{\rm s}}\left(H_{{\rm a,dem}}^{{\rm{\rm tube}}}\cos^{2}\phi-\pi H_{{\rm D,dem}}^{{\rm tube}}Q\cos\phi\right),$ (5)
where $H_{{\rm a,dem}}^{{\rm{\rm tube}}}$ and $H_{{\rm D,dem}}^{{\rm tube}}$ represent the anisotropy-like demagnetizing field within the DW [73] and DMI-like field which is specific to radially magnetized tube [51], respectively. These fields are expressed as
$H_{{\rm a,dem}}^{{\rm{\rm tube}}}=\frac{4\pi M_{{\rm s}}}{\rho_{1}^{2}-\rho_{0}^{2}}\int_{-\infty}^{+\infty}{\rm d}x\int_{0}^{+\infty}{\rm d}y\,\frac{\tanh x\left(\rho_{1}^{2}I_{11}^{-1}\left(\rho_{1},\rho_{1},y\right)-2\rho_{1}\rho_{0}I_{11}^{-1}\left(\rho_{1},\rho_{0},y\right)+\rho_{0}^{2}I_{11}^{-1}\left(\rho_{0},\rho_{0},y\right)\right)}{\cosh x\cosh\left(y-x\right)}\,,$ (6)
$H_{{\rm D,dem}}^{{\rm tube}}=\frac{8M_{{\rm s}}}{\rho_{1}^{2}-\rho_{0}^{2}}\int_{-\infty}^{+\infty}{\rm d}x\int_{0}^{+\infty}{\rm d}y\int_{\rho_{0}}^{\rho_{1}}{\rm d}\rho\,\frac{\tanh x\left(\rho_{1}I_{11}^{1}\left(\rho_{1},\rho,y\right)-\rho_{0}I_{11}^{1}\left(\rho_{0},\rho,y\right)\right)\rho}{\cosh\left(y-x\right)},$ (7)
@@
Figure 3: Energy densities per cross section in PMA nanotubes as a function of magnetization angle. The following parameters are used; $A=1\,{\rm uerg/cm}$ and $K_{\rm u}=8.3\,{\rm Merg/cm}^{3}$.
@@
The second term of Eq. (5) arises from the stray field from the magnetic domains in the radially magnetized tube (Fig. 4(b)). When the DW magnetization is aligned with the stray field, corresponding to the right-handed Néel DW, the energy of this configuration is at a minimum. This term consists of not only the interaction between geometrical magnetostatic charges but also the interaction between geometrical and surface magnetostatic charges [51]. The fundamental basis of this field lies in the asymmetry in the distribution of magnetic charges on the front and back surfaces of the magnetic layer. It should be noted that this field is specific to radially magnetized tubes and is not present in ribbons and other tubes with a different easy axis of magnetization.
@@
The exchange interaction increases with greater curvature, leading to a corresponding increase in $\left|H_{{\rm a,exc}}^{{\rm tube}}\right|$ as the radius decreases, shown in Fig. 5(a). Since this field is negative, Néel-type DWs are favored in systems with smaller radii. This characteristics is similar to the behavior observed in nanoribbons, where narrower wire width favor Néel-type DWs [73]. However, unlike in nanoribbons where this characteristics is attributed to exchange interactions, in nanotube, it is due to magnetostatic interactions, indicating that the origins of the two characteristics are different.
@@
The anisotropy-like demagnetization field from the magnetic charges in the DW, $H_{{\rm a,dem}}^{{\rm tube}}$, is primarily determined by the DW width and is thus independent of the tube radius, as shown in Fig. 5(b). This characteristics differs from the behavior observed in nanoribbons, where the anisotropy-like demagnetization field depends on the wire width [73]. In nanoribbons, as the wire width decreases, this effective field decreases and eventually reverses its sign at a certain width, favoring Néel-type DWs over Bloch-type DWs. This discrepancy arises from the presence of edges in nanowires. The absence of edges in tubes prevents the emergence of magnetic charges within Bloch-type DWs, making them preferred ($H_{{\rm a,dem}}^{{\rm tube}}>0$).
@@
In the radially magnetized tube, the total anisotropy and DMI fields acting on a DW are written by $H_{{\rm a}}=H_{{\rm a,exc}}^{{\rm tube}}+H_{{\rm a,dem}}^{{\rm tube}}$ and $H_{{\rm D}}=H_{{\rm D,dem}}^{{\rm tube}}+H_{{\rm DMI}}$, respectively. The magnetization angle of the DW in the equilibrium states can be expressed by the following equation [75]:
$\phi_{{\rm eq}}=\begin{cases}0&\frac{2H_{{\rm a}}}{\left|\pi H_{{\rm D}}\right|}<1\cap H_{{\rm D}}>0\\ \pi&\frac{2H_{{\rm a}}}{\left|\pi H_{{\rm D}}\right|}<1\cap H_{{\rm D}}<0\\ {\rm arccos}\left(\frac{\pi H_{{\rm D}}}{2H_{{\rm a}}}\right)&\frac{2H_{{\rm a}}}{\left|\pi H_{{\rm D}}\right|}>1\end{cases},$ (8)
@@
In nanoribbons, $H_{{\rm a}}$ arises exclusively from the anisotropy-like demagnetization field arising from magnetic charges within the DW, while $H_{{\rm D}}$ does not include any geometry-dependent terms. Consequently, in nanoribbons, the total anisotropy field is influenced by the nanoribbon width [73], whereas the total DMI field remains invariant to changes in dimensions.
Figure 6: Magnetization angle of DW in equilibrium states of PMA nanotubes. The following parameters are used; $A=1\,{\rm uerg/cm}$, $K_{\rm u}=8.3\,{\rm Merg/cm}^{3}$, and $t=1\,{\rm nm}$.
IV Spin-orbit torque-driven domain wall motion

### v2

# Spin-orbit torque driven motion of chiral domain walls induced by radial magnetization in nanotube geometry

@@ Abstract
We theoretically investigate chiral domain walls (DWs) formed in radially magnetized nanotubes composed of ultrathin layers with perpendicular magnetic anisotropy (PMA). Unlike those with in-plane magnetic anisotropy, the stable configurations of DWs in PMA nanotubes are influenced not only by exchange interactions but also by magnetostatic interactions induced by the radial component of magnetization. Particularly, the magnetostatic interactions lead to Dzyaloshinskii-Moriya interaction (DMI)-like effects that stabilize chiral Néel-type DWs. We derive expressions for the effective magnetic fields acting on DWs within PMA nanotubes and quantify spin-orbit torque (SOT) driven DW motion using an analytical one-dimensional model, which is validated by micromagnetic simulations. Our results show that the DMI-like field due to magnetostatic interactions can be as significant as the contribution of material-induced DMI in nanotubes with diameters below 100100\,nm. This implies that the direction and speed of DW motion in the PMA nanotubes could differ from those observed in flat nanoribbons composed of the same material. Furthermore, we demonstrate that DW velocity can be effectively controlled by adjusting the tube diameter and exchange stiffness constant of the magnetic layer, rather than relying solely on material-induced DMI. These insights are expected to greatly expand the potential applications of PMA nanotube-based DW devices.
@@ I Introduction
In this study, we focus on nanotubes as a promising structure for future 3D magnetic DW memory (DWM) [14, 15, 16, 17]. Magnetic nanotubes and nanocylinders can be fabricated using methods such as rolling up of thin films [18, 19, 20, 21], electroplating [22, 23, 24, 25, 26, 27, 28, 16, 29, 30, 31, 32, 33], and atomic layer deposition (ALD) [34, 35, 36, 37]. Compared to nanocylinders, nanotubes have the advantage of supplying higher current densities for the same applied current due to their smaller cross-sectional area, which is crucial for generating sufficient torque for DW motion. Although ALD techniques for magnetic structures are still in their early stages, recent observations of current-induced DW motion in ALD-Co layers [38] lead us to believe that magnetic nanotubes utilizing ALD techniques represent a promising candidate for the essential components of 3D-DWM.
@@
Previous theoretical research on the properties of DWs in nanotubes has primarily focused on systems with in-plane magnetic anisotropy (IMA) [39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50]. These studies have predominantly considered DW motion driven by conventional spin-transfer torque (STT) and Ørsted field [43, 50]. In contrast, this study aims to establish a theoretical framework for DW dynamics in radially magnetized nanotubes composed of thin films with perpendicular magnetic anisotropy (PMA) [21, 51]. Such systems are expected to utilize spin-orbit torque (SOT) [52, 53, 54], potentially enabling more efficient DW motion compared to traditional driving forces. Therefore, PMA nanotube-based DW shift registers are considered highly promising candidates for low power consumption data storage applications.
@@
Generally, geometrical effects are known to arise from exchange interactions [55, 56]; however, it has been predicted that geometrical effects due to magnetostatic interactions can also occur, depending on the curvature characteristics and the easy axis of magnetization [57]. Previous studies indicate that such magnetostatic interactions in tubular structures induce geometrical effects only in systems exhibiting a magnetization component along the radial axis of the tube [57, 58], leading to energy contributions similar to the Dzyaloshinskii-Moriya interaction (DMI) [59, 60]. This suggests the potential induction of chirality in DWs within tubular structures. The magnitude of DMI is crucial for the efficiency of SOT-driven DW motion [52], making it essential to theoretically predict the impact of these geometrical effects, to evaluate the potential applications of PMA nanotubes.
@@
The purpose of this study is to elucidate the characteristics of chiral DWs in PMA nanotubes and quantitatively predict the influence of geometrical effects on SOT-driven DW motion. Specifically, we derive expressions for the effective magnetic fields acting on DWs due to magnetostatic interactions and formulate an analytical model for SOT-driven DW motion. The validity of this model is confirmed through micromagnetic (µM) simulations.
@@
The remainder of this paper is organized as follows. In Section II, we describe the model of PMA nanotubes. Section III presents the expressions for effective magnetic fields due to geometrical effects in PMA nanotubes and discusses the characteristics of these fields and the stability of DWs as a function of nanotube diameter. In Section IV, we present the results of SOT-driven DW motion calculations, comparing the outcomes from the analytical one-dimensional model (1DM) [61, 62, 52, 63, 64] and µM simulations [63]. We also discuss the comparative shift characteristics of nanotubes and flat nanoribbons, as well as the impact of variations in tube diameter and magnetic properties on these characteristics. Finally, Section V summarizes the key findings and implications of this study.
@@ II Model of radially magnetized nanotube
Figure 1: Schematic illustration of a radially magnetized nanotube containing a DW. The magnetic structure for an up-down DW configuration ($Q=+1$) is presented. The enlarged diagram shows the magnetization configuration of the DW as seen from the inside of the tube.
The total energy of our model comprises contributions from the PMA, DMI, exchange interactions, and magnetostatic interactions (demagnetization) terms. We assume nanotubes made of typical magnetic materials consisting of 3dd transition metal alloys, characterized by the following parameters: the saturation magnetization $M_{{\rm s}}$ ($10^{6}\,{\rm A/m}$), the exchange stiffness constant $A$, the magnetic anisotropy constant perpendicular to the magnetic layer $K_{{\rm u}}$, and the DMI constant $D$. The sign of $D$ is defined under the assumption that the outer side of the tube is the substrate side. When $D>0$ ($D<0$), a right (left)-handed Néel-type DW is favored. To focus on the geometrical effects induced by magnetization configurations constrained by the nanotube shape and the magnetic anisotropy, the modulation of electronic states by the curvature of the nanotube [77] are neglected in our model ; instead, we assume these magnetic parameters to be constants.
In PMA nanotubes, the out-of-plane direction perpendicular to the film surface coincides with the radial direction of the tube, thus the PMA energy density can be expressed as $\varepsilon_{{\rm PMA}}=-K_{{\rm u}}m_{r}^{2}$. Additionally, the DMI and exchange energy densities can be respectively expressed as $\varepsilon_{{\rm DMI}}=D\left[m_{r}\frac{\partial}{\partial z}m_{z}-m_{z}\frac{\partial}{\partial z}m_{r}+\frac{1}{r}\left(m_{r}\frac{\partial}{\partial\varphi}m_{\varphi}-m_{\varphi}\frac{\partial}{\partial\varphi}m_{r}\right)\right]$ and $\varepsilon_{{\rm exc}}=A\sum_{i=x,y,z}\left[\left(\frac{\partial m_{i}}{\partial r}\right)^{2}+\frac{1}{r^{2}}\left(\frac{\partial m_{i}}{\partial\varphi}\right)^{2}+\left(\frac{\partial m_{i}}{\partial z}\right)^{2}\right],$ where $m_{x}=m_{r}\cos\varphi-m_{\varphi}\sin\varphi$ and $m_{y}=m_{r}\sin\varphi+m_{\varphi}\cos\varphi$. The energy of magnetistatic interaction is expressed as:
@@
$E_{{\rm dem,bb}}$ $=\frac{1}{2}\frac{\mu_{0}}{4\pi}\int{\rm d}^{3}\bm{x}\int{\rm d}^{3}\bm{x}^{\prime}\frac{\rho_{{\rm m}}\left(\bm{x}\right)\rho_{{\rm m}}\left(\bm{x}^{\prime}\right)}{\left|\bm{x}-\bm{x}^{\prime}\right|},$
$E_{{\rm dem,ss}}$ $=\frac{1}{2}\frac{\mu_{0}}{4\pi}\int{\rm d}^{2}\bm{x}\int{\rm d}^{2}\bm{x}^{\prime}\frac{\sigma_{{\rm m}}\left(\bm{x}\right)\sigma_{{\rm m}}\left(\bm{x}^{\prime}\right)}{\left|\bm{x}-\bm{x}^{\prime}\right|},$
$E_{{\rm dem,bs}}$ $=\frac{\mu_{0}}{4\pi}\int{\rm d}^{3}\bm{x}\int{\rm d}^{2}\bm{x}^{\prime}\frac{\rho_{{\rm m}}\left(\bm{x}\right)\sigma_{{\rm m}}\left(\bm{x}^{\prime}\right)}{\left|\bm{x}-\bm{x}^{\prime}\right|},$
where $\mu_{0}$ is the permeability of free space, $\sigma_{{\rm m}}\left(\bm{x}\right)=M_{\rm s}\bm{m}\left(\bm{x}\right)\cdot\bm{n}_{\perp}\left(\bm{x}\right)$ is the surface magnetic charge density, with $\bm{n}_{\perp}\left(\bm{x}\right)$ being the normal vector of the tube surface, and $\rho_{{\rm m}}\left(\bm{x}\right)=-M_{\rm s}\nabla\cdot\bm{m}\left(\bm{x}\right)$ is the volume magnetic charge density [78].
Figure 2: Phase diagram of magnetic domain structures in the PMA nanotube. The dashed line represents the boundary of stable states as determined by Eq. (2). The markers indicate conditions validated through µM simulations. The following parameters are used; $A=10^{-11}\,{\rm J/m}$ and $t=1\,{\rm nm}$.
Assuming the demagnetizing factor in the radial direction is equal to 1 due to $t\ll R_{0}$, and comparing the energy density of the uniform radial magnetization structure with that of the uniform axial magnetization structure, the effective magnetic anisotropy constant for PMA layers of nanotubes is given by
$K_{{\rm eff}}=K_{{\rm u}}-\frac{\mu_{0}}{2}M_{{\rm s}}^{2}+\frac{A}{\left(R_{1}-t/2\right)t}\ln\frac{R_{1}}{R_{1}-t},$ (2)
where the second term represents the demagnetization energy density, and the third term accounts for the exchange energy density associated with the radial domain state [78], which is specific to tubular structures. Figure 2 shows the phase diagram of magnetic domain structures in the PMA nanotube, validated through µM simulations (see Appendix A). When $K_{{\rm eff}}>0$, the easy axes of magnetization are oriented radially; otherwise, the easy axis aligns along the axial direction.
@@ III Characteristics of domain walls
In this section, we derive the energy densities per nanotube cross-sectional area and present expressions for the effective fields acting on the DWs in PMA nanotubes.
Hereafter, we discuss structures where the thickness $t$ of the magnetic layer in the nanotube is smaller than the exchange length $\lambda=\sqrt{A/K_{{\rm eff}}}$. The lack of edges in the nanotube’s cross-section, which is perpendicular to the nanotube’s axis, leads to high symmetry, making it reasonable to assume uniform magnetization across the cross-section when $t<\lambda$. Therefore, the magnetization structure at $z=X$ with the magnetization angle of $\phi$ is expressed as $m_{r}=-Q\,{\tanh\frac{z-X}{\lambda}}$, $m_{\varphi}={\rm sech}{\frac{z-X}{\lambda}\sin\phi}$, and $m_{z}={\rm sech}{\frac{z-X}{\lambda}}\cos\phi$, where the DW half-width is approximately equal to the exchange length $\lambda$ and $Q$ represents the DW configuration, defined as $Q=+1$ for an up-down ($\uparrow\downarrow$) wall (see Fig. 1) and Q=1Q=-1 for a down-up ($\downarrow\uparrow$) wall [61, 62, 52, 63, 64].
The PMA, DMI, and exchange energy densities per nanotube cross-sectional area are expressed as $w_{{\rm PMA,DMI,exc}}=\int{\rm d}z\,\varepsilon_{{\rm PMA,DMI,exc}}$. The integral result of the PMA term does not depend on $\phi$ and thus does not affect the DW dynamics (see Eq. (9)).
@@
$w_{{\rm DMI}}\left(\phi\right)=-\pi\mu_{0}M_{{\rm s}}\lambda H_{{\rm DMI}}Q\cos\phi,\\$ (3)
where $H_{{\rm DMI}}=\frac{D}{\mu_{0}M_{{\rm s}}\lambda}$ is the material-induced DMI field [79, 63, 80].
The integral result of the $\phi$-dependent term of exchange interaction is given by
$w_{{\rm exc}}\left(\phi\right)=\mu_{0}M_{{\rm s}}\lambda H_{{\rm a,exc}}^{{\rm tube}}\cos^{2}\phi,\\$ (4)
where $H_{{\rm a,exc}}^{{\rm tube}}=-\frac{4A}{\mu_{0}M_{{\rm s}}}\frac{1}{R_{1}^{2}-R_{0}^{2}}\ln\frac{R_{1}}{R_{0}}$ is the curvilinear-geometry-induced anisotropy-like field [55, 56]. This result is obtained by integrating $\varepsilon_{{\rm exc}}=\frac{A}{r^{2}}\left(1+\frac{r^{2}}{\lambda^{2}}{\rm sech}^{2}\frac{z-X}{\lambda}-{\rm sech}^{2}\frac{z-X}{\lambda}\cos^{2}\phi\right)$ with respect to $X$ [81], and is consistent with the result for IMA tubes with an azimuthal easy axis of magnetization [50]. In the case where $H_{{\rm a,exc}}^{{\rm tube}}$ is positive (negative), the Bloch (Néel)-type DW is favored.
The energy density of the $\phi$-dependent term of magnetostatic interaction is given by the following equation, as derived in Appendix B:
$w_{{\rm dem}}\left(\phi\right)=\mu_{0}M_{{\rm s}}\lambda\left(H_{{\rm a,dem}}^{{\rm{\rm tube}}}\cos^{2}\phi-\pi QH_{{\rm D,dem}}^{{\rm tube}}\cos\phi\right),$ (5)
where $H_{{\rm a,dem}}^{{\rm{\rm tube}}}$ and $H_{{\rm D,dem}}^{{\rm tube}}$ represent the anisotropy-like demagnetizing field within the DW [82] and DMI-like field which is specific to radially magnetized tubes [57], respectively. These fields are expressed as
$H_{{\rm a,dem}}^{{\rm{\rm tube}}}$ $=\frac{2M_{{\rm s}}}{\rho_{1}^{2}-\rho_{0}^{2}}\int_{0}^{+\infty}{\rm d}y\,\frac{1}{\sinh y}\left(\frac{y}{\tanh y}-1\right)$
$\quad\times\Bigl{(}\rho_{1}^{2}I_{11}^{-1}\left(\rho_{1},\rho_{1},y\right)-2\rho_{1}\rho_{0}I_{11}^{-1}\left(\rho_{1},\rho_{0},y\right)$
$\quad\quad+\rho_{0}^{2}I_{11}^{-1}\left(\rho_{0},\rho_{0},y\right)\Bigr{)},$ (6)
$H_{{\rm D,dem}}^{{\rm tube}}$ $=\frac{2M_{{\rm s}}}{\rho_{1}^{2}-\rho_{0}^{2}}\int_{0}^{+\infty}{\rm d}y\,\tanh\frac{y}{2}\int_{\rho_{0}}^{\rho_{1}}{\rm d}\rho\,\rho$
$\quad\times\left(\rho_{1}I_{11}^{1}\left(\rho_{1},\rho,y\right)-\rho_{0}I_{11}^{1}\left(\rho_{0},\rho,y\right)\right),$ (7)
@@
Figure 3: Energy densities per cross-sectional area in PMA nanotubes as a function of magnetization angle. The following parameters are used; $A=10^{-11}\,{\rm J/m}$ and $K_{\rm u}=0.83\times 10^{6}\,{\rm J/m^{3}}$.
@@
The second term of Eq. (5) arises from the stray field from the magnetic domains in the radially magnetized tube (Fig. 4(b)). When the DW magnetization is aligned with the stray field, corresponding to the right-handed Néel-type DW, the energy of this configuration is at a minimum. This term consists of not only the interaction between geometrical magnetostatic charges but also the interaction between geometrical and surface magnetostatic charges [57]. The fundamental basis of this field lies in the asymmetry in the distribution of magnetic charges on the inner and outer surfaces of the nanotube. It should be noted that this field is specific to radially magnetized tubes and is not present in flat ribbons and other tubes with a different easy axis of magnetization.
@@
The exchange interaction increases with greater curvature, leading to a corresponding increase in $\left|H_{{\rm a,exc}}^{{\rm tube}}\right|$ as the radius decreases, shown in Fig. 5(a). Since this field is negative, Néel-type DWs are favored in systems with smaller radii. This characteristic is similar to the behavior observed in flat nanoribbons, where narrower wire width favor Néel-type DWs [82]. However, unlike in flat nanoribbons where this characteristic is attributed to magnetostatic interactions, in nanotubes, it is due to exchange interactions, indicating that the origins of the two characteristics are different.
@@
The anisotropy-like demagnetizing field from the magnetic charges in the DW, $H_{{\rm a,dem}}^{{\rm tube}}$, is primarily determined by the DW width and is thus independent of the tube radius, as shown in Fig. 5(b). This characteristic differs from the behavior observed in flat nanoribbons, where the anisotropy-like demagnetizing field depends on the wire width [82]. In flat nanoribbons, as the wire width decreases, this effective field decreases and eventually reverses its sign at a certain width, favoring Néel-type DWs over Bloch-type DWs. This discrepancy arises from the presence of edges in nanowires. The lack of edges in the tube’s cross-section prevents the emergence of magnetic charges within Bloch-type DWs, making them preferred ($H_{{\rm a,dem}}^{{\rm tube}}>0$).
@@
In PMA nanotubes, the total anisotropy and DMI fields acting on a DW are written by $H_{{\rm a}}=H_{{\rm a,exc}}^{{\rm tube}}+H_{{\rm a,dem}}^{{\rm tube}}$ and $H_{{\rm D}}=H_{{\rm D,dem}}^{{\rm tube}}+H_{{\rm DMI}}$, respectively. The magnetization angle of the DW in the equilibrium states can be expressed by the following equation [84]:
@@
where we assume that $Q=+1$ and consider only the range of $\phi$ from 0 to $\pi$ due to the symmetry. The equilibrium angle is influenced by both the material-induced DMI constant and the radius of the nanotube, as shown in Fig. 6. As the radius decreases, the intermediate region between the Bloch- and Néel-type DWs contracts, reflecting the radius dependence of $H_{{\rm a,exc}}^{{\rm tube}}$. Additionally, the boundary between the left- and right-handed Néel configurations shift toward the negative side of $D$ as the radius decreases, reflecting the radius dependence of $H_{{\rm D,dem}}^{{\rm tube}}$. For $R_{1}\lesssim 25\,{\rm nm}$, no intermediate states exist between the Bloch- and Néel-type DWs, leading to a discontinuous transition between left- and right-handed Néel configurations. This discontinuity is due to the absence of an effective field to stabilize a Bloch-type DW, as $H_{\rm a}$ becomes negative for $R_{1}\lesssim 25\,{\rm nm}$. This corresponds to the condition $2H_{\rm a}/\left|\pi H_{\rm D}\right|<1$ in Eq. (8). In the PMA nanotube composed of typical 3dd transition metal alloys (with $M_{\rm s}=10^{6}\,{\rm A/m}$ assumed in this model), it is confirmed that geometrical effects contribute on the order of $0.1\,{\rm mJ/m^{2}}$, comparable to the material-induced DMI [71, 80].
In flat nanoribbons, $H_{{\rm a}}$ arises exclusively from the anisotropy-like demagnetizing field arising from magnetic charges within the DW, while $H_{{\rm D}}$ does not include any geometry-dependent terms. Consequently, in flat nanoribbons, the total anisotropy field is influenced by the nanoribbon width [82], whereas the total DMI field remains invariant to changes in dimensions.
Figure 6: Magnetization angle of DW in equilibrium states of PMA nanotubes. The following parameters are used; $A=10^{-11}\,{\rm J/m}$, $K_{\rm u}=0.83\times 10^{6}\,{\rm J/m^{3}}$, and $t=1\,{\rm nm}$.
IV Spin-orbit torque driven domain wall motion